# Beyond Dialogue: A Profile-Dialogue Alignment Framework Towards General Role-Playing Language Model

### v3

@@ Appendix C Human Verification of Evaluation
In the Auto-Evaluation part, GPT-4o evaluated role-playing dialogues generated by different models across eight dimensions: Character, Style, Emotion, Relationship, Personality, Human-likeness, Role Choice, and Coherence. The human verification process involved two key metrics: Validity and Cosine similarity. Validity assesses the soundness of GPT-4o’s evaluation justifications, while Cosine similarity measures the agreement between GPT-4o’s evaluation results and those of human evaluators.
@@
Validity 1.0 ± 0.0 1.0 ± 0.0 0.94 ± 0.02 1.0 ± 0.0 1.0 ± 0.0 0.89 ± 0.04 0.9 ± 0.03 0.92 ± 0.03 0.98 ± 0.01 0.93 ± 0.03 0.88 ± 0.03 0.92 ± 0.03 0.99 ± 0.01
Cosine - 0.96 ± 0.01 0.94 ± 0.01 0.94 ± 0.01 0.99 ± 0.0 0.98 ± 0.01 0.78 ± 0.0 0.91 ± 0.03 0.99 ± 0.0
Table 9: Human Verification of GPT-4o’s Evaluation in Auto-Generation and Auto-Test. Validity reflects the percentage of evaluations deemed reasonable by human evaluators. Cosine similarity measures the agreement between GPT-4o’s evaluation results and those of human evaluators.
@@ Appendix D Implementation Details
Training. We fully finetuning all models for 3 epochs with 833 steps with the 128 global batch size. We set the training sequence length to 4,096. The learning rate is 3e-5. The + RPA & CC & CSERP represents models were trained using Align Role-Playing dialogue with profiles, Chit-Chat, and CSERP data in a ratio of 1:4:5, while + RP & CC represents a ratio of 1:9 using non-aligned dialogue data. In both cases, the proportion of English to Chinese data within the Chit-Chat data is consistent with that in the Role-Playing data.
@@
Hardware and System Configuration We use 8 NVIDIA RTX A100 GPUs with 1TB of memory for training and a single A100 GPU for inference. The system operates on Linux version 4.14.105-1-tlinux3-0013.
@@ Appendix E Cost Analysis
Role-playing Dialogue Dataset Construction
In this section, we calculated the costs associated with constructing the role-playing dialogue dataset, ensuring both cost-effectiveness and efficiency.
Automated Pipeline. We employed GPT-4o for dialogue extraction and reconstruction, with an average cost of \$0.016 per item. Initial chunk checking and scenario construction were handled by open-source models, and excluded from the cost analysis. The comprehensive extraction and verification process using GPT-4o amounted to \$0.048 per item.
Human Annotation. For comparison, human annotation costs were \$16/hour (approximately \$1 per item), significantly higher than the automated method. Human annotation not only incurs over 20 times the cost per item but also is substantially slower.
Our findings highlight that the pipeline we employed for constructing the role-playing dialogue dataset is both low-cost and highly efficient. Despite the significantly lower cost, we ensured the reliability of the automated process by conducting a thorough human review of the dialogues extracted by GPT-4o. As detailed in §A, the accuracy of GPT-4o’s outputs was found to exceed 96%, demonstrating that the constructed dataset is trustworthy and of high quality. In conclusion, our automated pipeline not only reduces costs but also accelerates the process of dataset construction, providing a scalable solution for future projects while maintaining a high standard of data accuracy.
Alignment and Evaluation
In this section, we present a detailed cost analysis of the alignment and evaluation process for Beyond Dialogue framework. Leveraging our automated pipeline, we conducted a comprehensive evaluation across multiple dimensions, including Character, Style, Emotion, Relationship, and Personality, as well as additional tasks such as Human-likeness, Coherence, and Role Choice. The key components of our cost analysis are outlined below.
Automated Alignment. The alignment of data across these five dimensions (Character, Style, Emotion, Relationship, and Personality) was performed using our automated pipeline, with an average cost of \$0.012 per data item. In contrast, manual annotation for alignment would cost approximately \$16 per hour, which translates to about \$0.5 per data item. This demonstrates that our automated method reduces the cost by a factor of over 40 compared to traditional human annotation.
Automated Evaluation. Before the evaluation, our automated process includes an initial cost for generating dialogues with the model. Each model interaction involves a 5-round dialogue, costing approximately \$0.02 per data item. Following this, our evaluation process expands beyond the alignment dimensions to include three additional task(Human-likeness, Coherence, and Role Choice), which are essential for ensuring that the dialogues not only align with predefined profiles but also exhibit realism and logical consistency. The cost of evaluating each data item across these eight dimensions, including the initial model interaction, is approximately \$0.1 per data item. For comparison, manual evaluation, which involves interacting with the model and performing the eight evaluation tasks, takes about 12 minutes per data item, resulting in a cost of approximately \$5 per item. This comparison highlights that our automated evaluation process is over 50 times more cost-effective than traditional human annotation.
This evaluation process, which includes both automated alignment and additional assessment dimensions, underscores the efficiency and cost-effectiveness of our pipeline. By leveraging automation, we achieve a comprehensive evaluation at a fraction of the cost and time required for manual annotation, while still maintaining high standards of accuracy and quality, as demonstrated in §C. In conclusion, our alignment and evaluation framework not only ensures that the constructed role-playing dialogue dataset is aligned with the intended profiles but also that it is evaluated rigorously for quality and coherence, all while significantly reducing costs compared to traditional methods.
@@ Appendix F Prompt Templates
Dataset Construction Prompts
@@
    Emotion Alignment: This prompt requires the analysis of Ekman basic emotions (Ekman 1992) from the dialogue of a specified role, assessing emotional expressions in context to the scene (Table 17).
@@
Auto Dialogue and Evaluation Prompts
The prompt templates used in the automated dialogue pipeline are as follows: chat role generation (Table 20), dialogue scenario generation (Table 21), and generation of Emotion and Relationship (Tables 22 and 23). Additionally, prompts for playing the chat role and prompts for playing the evaluated roles are presented in Tables 24 and 25.
In our role-playing training data framework, evaluation prompts aim to assess dialogues for adherence to established profiles, reflecting the quality of role portrayal. To this end, evaluation prompts for the CSERP dimensions (Character, Style, Emotion, Relationship, Personality) align with the templates in Alignment Prompts in §F.
@@ Appendix G Case Study
In this section, we sample several cases from the role-playing dialogue and alignment tasks to illustrate the effectiveness of our proposed framework.
Role-playing Dialogue
In this subsection, we present a series of dialogue cases in both Chinese and English, with the original Chinese dialogues translated into English. The cases are illustrated in Figures 11 to 18, showcasing the role-playing capabilities of the Qwen2-7B and Mistral-Nemo models.
Through our case studies, we observe that the models, after training, exhibit significant improvements in adhering to the predefined role profiles and producing more human-like responses. For instance, the trained Mistral-Nemo model, when role-playing as Hamlet, consistently generates responses that are polite and eloquent, reflecting the role’s sophisticated nature. Similarly, the trained Qwen2-7B model, when portraying Zhou Botong in Chinese, captures the essence of his “Old Urchin” persona, demonstrating a mischievous and playful demeanor in its output. These results underscore the effectiveness of our training approach in enhancing the models’ ability to represent complex role traits in role-playing dialogues accurately.
Alignment Task
This subsection delves into the alignment task within the CSERP framework, which encompasses five critical dimensions: Character, Style, Emotion, Relationship, and Personality. For Character, Style, and Personality (CSP), alignment is assessed through word recall, while Emotion and Relationship (EP) are evaluated using a 0-10 scoring system. Thus, we selected one task from each of the aforementioned types for case demonstration, specifically focusing on Character and Emotion, as shown in Figures 19 to 22.
The untrained Qwen2-7B model struggles with fine-grained alignment of profiles and dialogues, often failing to follow instructions effectively. After training, both Mistral-Nemo and Qwen2-7B models perform closer to GPT-4o.
@@
You are an expert with a deep understanding of literary works, skilled at analyzing and extracting the core elements of literature. Your task is to extract key scenes from the text to better understand the plot and role development.
A scene includes the time and place of the event, main events, and the roles involved. Do not include role dialogues in the scene. Here is an example of a provided scene:
@@
Now, based on the above requirements, extract the key scene from the text and describe it accordingly. Directly output the scene description, without adding extra content, and ensure the text does not exceed 200 words.
@@
You are an expert with a strong background in literature and psychology, skilled at analyzing and interpreting the role traits and dialogue performances of roles from texts. Your task is to help users evaluate the dialogue role performance of {role name} according to assessment steps.
@@
5. Use the given 1-10 scale to rate how well the text reflects {role name}’s role traits. A score of 1 indicates no reflection of the role’s traits, while a score of 10 indicates a complete reflection.
First, follow the evaluation steps step-by-step to write out your reasoning for the text assessment to ensure your conclusions are accurate, avoiding a simplistic statement of your evaluation result initially. Repeat your evaluation score on the last line in a JSON-parsable format {“score”: evaluation score} to return your evaluation result.
@@
Your goal is to extract structured information from the user’s input that matches the form described below. When extracting information please make sure it matches the type information exactly. Do not add any attributes that do not appear in the schema shown below.
@@
role: string // The role who is speaking or performing an action, use context to predict the name of the role. Use scene or narrator if no one speak
@@
action: string // The actions performed by the roles in the text, A high-level summary of a role’s behavior. action equals “dialogue”. if it’s no dialogue, summarize role’s behavior in sentence
@@
Do NOT add any clarifying information. Output MUST follow the schema above. Do NOT add any additional columns that do not appear in the schema.
@@
You are an expert in scene analysis, skilled at analyzing and extracting key information from texts. Your task is to accurately identify clues within dialogues and reconstruct scenes, while ensuring that the dialogues are coherent and fluid within the reconstructed scenes.
@@
Now, based on the above requirements, reconstruct the sub-scene where the dialogue takes place, and describe it accordingly. Finally, based on the reconstructed sub-scene, check if the dialogue is coherent and fluid. You may briefly analyze the scene context and its coherence, then return your evaluation result in a JSON-parsable format as follows:
{“scene”: “reconstructed scene description”, “coherence”: 1/0} Where “coherence” of 1 indicates the dialogue is coherent and fluid with the scene, and 0 indicates the dialogue is not coherent with the scene.
Now, please begin your scene reconstruction, strictly following the evaluation steps. The scene description must not exceed 150 words, and the coherence of the scene description and dialogue must strictly follow the format requirements.
@@
You are an expert in the fields of literature and psychology, skilled at analyzing and interpreting the role traits and dialogue performances in texts. Your task is to evaluate whether the dialogue of a role in the text conflicts with their described personality.
@@
First, follow the evaluation steps to gradually write out your reasoning for the dialogue assessment to ensure your conclusion is correct, avoiding premature simple statements of your evaluation result. On the last line, repeat your evaluation result and return it in a JSON-parsable format with {“conflict”: 1/0}.
@@
Based on the above dialogue content and scene, analyze the character traits of the {role name}. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
@@
{”character”: “trait1, trait2…”}
Now, please begin your analysis of {role name}’s character. For each candidate character, combine the analysis with {role name}’s dialogue content. Finally, select the character traits from the [Candidate Character Set]
@@
Based on the dialogue content and scene above, analyze the speaking style of the {role name}. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
@@
Now, please begin your analysis of {role name}’s speaking style. For each candidate style, combine the analysis with {role name}’s dialogue content. Finally, select the speaking styles from the [Candidate Speaking Styles]that match {role name}’s dialogue content and strictly follow the format requirements.
@@
Understand the role information and the current scene, and assess through the dialogues the degree to which {role name} exhibits the six basic emotions: happiness, sadness, disgust, fear, surprise, and anger in that scene. Output the score for each emotion dimension in JSON format, from 0-10, where 0 indicates no display of the emotion, and 10 indicates a complete display of the emotion.
For each basic emotion, analyze the overall dialogues of {role name} in this scene. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation results in a JSON-parsable format as follows:
@@
You are an emotional analysis expert, proficient in emotional analysis, psychology, dialogue understanding, and interpersonal relationship assessment. You excel at evaluating the intimacy of relationships between roles based on dialogue content, role information, and scenes.
@@
Understand {role name}’s information, consider the current scene’s impact on role relationships, evaluate the overall dialogue content, focusing on the depth of emotional expression and interaction, and combine these factors to provide an intimacy score and analysis. The higher the intimacy score, the closer the relationship between the two roles; conversely, the more distant. The intimacy score ranges from 0-10, where 0 represents the most distant relationships, indicating strangers, hostility, indifference, etc., and 10 represents the closest relationships, such as lovers, kin, or friends.
Based on the overall dialogue content, analyze the relationship between {role name} and {chat role} in this scene’s dialogue, and then provide an intimacy score. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation result in a JSON-parsable format as follows:
@@
The 8 letters of the MBTI correspond as follows: Introverted (I) / Extraverted (E); Intuitive (N) / Sensing (S); Thinking (T) / Feeling (F); Judging (J) / Perceiving (P).
@@
Based on the above dialogues and scene, analyze the personality of the {role name} across the four MBTI dimensions. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation result in a JSON-parsable format as follows:
@@
You need to design a new role description that will converse with {role name}, ensuring that the dialogue with this role effectively reflects {role name}’s personality, character traits, and speaking style.
@@
You need to creatively construct a new role setting to dialogue with {role name}, based on the traits of {role name}. This new role should not appear in any works related to {role name}. The description of the new role should include the role’s name and a brief personal description, to be output in JSON format like:
@@
You need to construct an engaging scene based on the information of roles A and B. The scene and roles’ actions must be consistent with the settings of both roles, ideally within 50-100 words, and consistent with the world they inhabit. The output should be in JSON format, like
@@
Now, please create a scene that fits the settings of both roles and is engaging. The scene should not directly include roles’ dialogues. Ensure that the output format meets the specified requirements.
@@
You need to assign six basic emotions to the {role name} in a specific scene: happiness, sadness, disgust, fear, surprise, and anger, based on the dialogue role’s information and the scene of the dialogue.
@@
Understand the role descriptions and the current scene, and assign the six basic emotions reflected in {role name}’s statements in that scene: happiness, sadness, disgust, fear, surprise, and anger. Output the score for each emotion dimension in JSON format, from 0-10, where 0 means the emotion is not displayed at all, and 10 means the emotion is fully displayed.
Please analyze in a few brief sentences the scores for the six basic emotions that {role name} should exhibit in this scene, avoiding a simplistic statement of your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation results in a JSON-parsable format as follows:
@@
Now, strictly follow the requirements to analyze the scene and role information, and assign emotional scores to {role name} that must be consistent with the role’s settings and the current scene. The analysis should be brief, not too lengthy, and avoid additional content. The final emotion scores must strictly follow the format requirements.
@@
You are an emotional analysis expert, proficient in psychology and interpersonal relationship assessment, skilled at initializing the intimacy level of relationships based on roles’ personalities and scenarios. You need to analyze the intimacy level between the {role name} and {chat role} by analyzing the information of both roles and the content of the scene.
@@
Understand {role name}’s personality and consider the current scene’s impact on the roles’ relationships, such as environment and context. Integrate the above factors to initialize their intimacy score. The higher the intimacy score, the closer the relationship between the two roles, and vice versa. The intimacy score ranges from 0-10, where 0 represents the most distant relationships, which can indicate strangers, hostility, indifference, etc., and 10 represents the closest relationships, which can include lovers, family, and friends.
Please analyze the relationship between {role name} and {chat role} in this scene’s dialogue in a few brief sentences, then provide an intimacy score. Avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation result in a JSON-parsable format as follows:

### v6

@@ Appendix C Human Verification of Evaluation
C.1 Human Annotation Process
For our human annotation process, participants were recruited from Mainland China, where they were employed by a company that specializes in data annotation. All participants were proficient in English, ensuring that the annotation tasks were completed accurately and efficiently. The compensation for these annotators was set at \$16 per hour, which corresponds to roughly \$1 per item annotated. This pay rate was determined based on the expected complexity of the tasks and the standard industry rates for such work in China.
Given the demographic characteristics of our participants (Chinese employees with strong English skills), this wage is considered fair within the local context. The annotators were able to work at a pace that aligned with this compensation, and the hourly rate was calibrated to account for the expected time it would take to complete each task. The average task duration was calculated through pilot testing, which helped us determine that \$16 per hour was appropriate for the time investment required.
C.2 Human Verification Results
@@
In the Auto-Evaluation part, GPT-4o evaluated role-playing dialogues generated by different models across eight dimensions: Character, Style, Emotion, Relationship, Personality, Human-likeness, Role Choice, and Coherence. The human verification process involved two key metrics: Validity and Similarity. Validity assesses the soundness of GPT-4o’s evaluation justifications, while Similarity measures the agreement between GPT-4o’s evaluation results and those of human evaluators.
@@
Validity 1.0 1.0 0.94 1.0 1.0 0.89 0.9 0.92 0.98 0.93 0.88 0.92 0.99
[…]
$\bm{\kappa}$ 0.85 0.77 N/A N/A 0.92 0.74 0.88 N/A
$\bm{\alpha}$ N/A 0.85 0.77 0.83 0.87 0.92 0.74 0.88 N/A
Acc 0.94 0.93 N/A N/A 0.97 0.73 0.96 0.99
F1 0.94 0.93 N/A N/A 0.97 0.72 0.95 0.99
Jaccard 0.90 0.89 N/A N/A 0.95 0.70 0.92 0.98
Table 12: Human Verification of GPT-4o’s Evaluation in Auto-Generation and Auto-Test. Validity reflects the percentage of evaluations deemed reasonable by human evaluators. Cosine similarity, Cohen’s Kappa ($\bm{\kappa}$), Krippendorff’s Alpha ($\bm{\alpha}$), Accuracy (Acc), F1 Score, and Jaccard Similarity (Jaccard) are used to assess the agreement between GPT-4o’s evaluation results and those of human evaluators. N/A indicates that the metric is not applicable for that specific evaluation.
To further assess the consistency and reliability of GPT-4o’s evaluations, additional inter-rater metrics were applied, including Cohen’s Kappa ($\kappa$), Krippendorff’s Alpha ($\alpha$), Accuracy (Acc), F1 Score, and Jaccard Similarity. These metrics show high consistency across most dimensions, particularly in Character, Style, Personality, and Role Choice, supporting the reliability of GPT-4o’s evaluation process and its alignment with human judgment.
@@ Appendix D Implementation Details
Training. We fully finetuning all models for 3 epochs with 833 steps with the 128 global batch size. We set the training sequence length to 4,096. The learning rate is 3e-5. The + RPA & CC & CSERP represents models were trained using Align Role-Playing dialogue with profiles, Chit-Chat, and CSERP data in a ratio of 1:4:5, while + RP & CC represents a ratio of 1:9 using non-aligned dialogue data. In both cases, the proportion of English to Chinese data within the Chit-Chat data is consistent with that in the Role-Playing data. In Ablation Study, each of the alignment tasks in $D_{a}$ was removed one by one. For each removed task, its corresponding training data volume was replaced by an equivalent amount of data from $D_{c}$, ensuring that the overall dataset size remained constant. The training settings for the ablation experiments were identical to those described above.
@@ Appendix E Cost Analysis
Process Automated Method Automated Cost (per item) Human Cost (per item)
Dialogue Extraction GPT-4o extraction and verification \$0.048 \$1 (\$16/hour)
Alignment (CSERP) Automated pipeline (Character, Style, Emotion, Relationship, Personality) \$0.012 \$0.5 (\$16/hour)
Model Interaction 5-round dialogue generation \$0.02 \$1 (\$16/hour)
Evaluation Automated evaluation (Human-likeness, Coherence, Role Choice, etc.) \$0.1 \$5 (\$16/hour, 12 min/item)
Total Cost per Item \$0.16 \$7.5
Table 13: Detailed Cost Analysis of Dataset Construction, Alignment, and Evaluation
As shown in Table 13, our automated pipeline reduces costs significantly compared to manual annotation. Dialogue extraction is 20× cheaper, alignment 40× cheaper, and evaluation 50× cheaper than human methods. Overall, the automated approach costs just \$0.16 per item, compared to \$7.5 for manual annotation, achieving a 47× cost reduction while maintaining high accuracy (96%, as verified in § A.2).
@@ Appendix F Prompt Templates
F.1 Dataset Construction Prompts
@@ F.2 Alignment Prompts
    Emotion Alignment: This prompt requires the analysis of Ekman basic emotions (Ekman, 1992) from the dialogue of a specified role, assessing emotional expressions in context to the scene (Table 21).
@@
F.3 Auto Dialogue and Evaluation Prompts
The prompt templates used in the automated dialogue pipeline are as follows: chat role generation (Table 24), dialogue scenario generation (Table 25), and generation of Emotion and Relationship (Tables 26 and 27). Additionally, prompts for playing the chat role and prompts for playing the evaluated roles are presented in Tables 28 and 29.
In our role-playing training data framework, evaluation prompts aim to assess dialogues for adherence to established profiles, reflecting the quality of role portrayal. To this end, evaluation prompts for the CSERP dimensions (Character, Style, Emotion, Relationship, Personality) align with the templates in Alignment Prompts in § F.2.
@@ Appendix G Case Study
In this section, we random sample several cases from the role-playing dialogue and alignment tasks in test data to illustrate the effectiveness of our proposed framework.
G.1 Role-playing Dialogue
In this subsection, we present a series of dialogue cases in both Chinese and English, with the original Chinese dialogues translated into English. The cases are illustrated in Figures 11 to 13, showcasing the role-playing capabilities of the Qwen2-7B and Mistral-Nemo models.
Through our case studies, we observe that the models, after training, exhibit significant improvements in adhering to the predefined role profiles and producing more human-like responses. For instance, the trained Mistral-Nemo model, when role-playing as Hamlet (Figure 12), consistently generates responses that are polite and eloquent, reflecting the role’s sophisticated nature. Similarly, the trained Qwen2-7B model, when portraying Zhou Botong in Chinese (Figure 13), captures the essence of his “Old Urchin” persona, demonstrating a mischievous and playful demeanor in its output. These results underscore the effectiveness of our training approach in enhancing the models’ ability to represent complex role traits in role-playing dialogues accurately.
G.2 Alignment Task
This subsection delves into the alignment task within the CSERP framework, which encompasses five critical dimensions: Character, Style, Emotion, Relationship, and Personality. For Character, Style, and Personality (CSP), alignment is assessed through word recall, while Emotion and Relationship (EP) are evaluated using a 0-10 scoring system. Thus, we selected one task from each of the aforementioned types for case demonstration, specifically focusing on Character and Emotion, as shown in Figures 14 to 17.
The alignment tasks achieve fine-grained alignment between the role profiles and dialogues at the sentence level by applying the Chain-of-Thought (COT) method during the data generation process. As shown in Figure 14, the sentence “It is only my own weakness. But tell me, how did you learn to mend such injuries?” spoken by “Dany” is aligned with the “Resilient” trait in her profile. This demonstrates how our training data establishes an explicit profile-to-dialogue sentence connection.
As shown in Figure 14, the untrained Qwen2-7B model struggles with fine-grained alignment of profiles and dialogues, often failing to follow instructions effectively. After training, both Mistral-Nemo and Qwen2-7B models perform closer to GPT-4o.
Appendix H Open Access and Licensing
The code used in this study is released under the Apache 2.0 License. The associated code repository is publicly available for use, modification, and distribution in compliance with the terms of the Apache 2.0 License.
The dataset used in this research is shared under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license. This dataset is available for non-commercial use and can be redistributed and modified under the terms specified by the license.
The code and dataset are provided in the supplementary files and will be made publicly available via open-source links upon acceptance of the paper. Detailed access instructions and relevant links will be included in the final version of the paper.
@@ Appendix H Open Access and Licensing
You are an expert with a deep understanding of literary works, skilled at analyzing and extracting the core elements of literature.
Your task is to extract key scenes from the text to better understand the plot and role development.
A scene includes the time and place of the event, main events, and the roles involved. Do not include role dialogues in the scene.
Here is an example of a provided scene:
@@
Now, based on the above requirements, extract the key scene from the text and describe it accordingly.
Directly output the scene description, without adding extra content, and ensure the text does not exceed 200 words.
@@
You are an expert with a strong background in literature and psychology, skilled at analyzing and interpreting the role traits and dialogue performances of roles from texts.
Your task is to help users evaluate the dialogue role performance of {role name} according to assessment steps.
@@
5. Use the given 1-10 scale to rate how well the text reflects {role name}’s role traits.
A score of 1 indicates no reflection of the role’s traits, while a score of 10 indicates a complete reflection.
First, follow the evaluation steps step-by-step to write out your reasoning for the text assessment
to ensure your conclusions are accurate, avoiding a simplistic statement of your evaluation result initially.
Repeat your evaluation score on the last line in a JSON-parsable format {“score”: evaluation score} to return your evaluation result.
@@
Your goal is to extract structured information from the user’s input that matches the form described below.
When extracting information please make sure it matches the type information exactly.
Do not add any attributes that do not appear in the schema shown below.
@@
role: string // The role who is speaking or performing an action, use context to predict the name of the role. Use ‘scene‘ or ‘narrator‘ if no one speak
@@
action: string // The actions performed by the roles in the text, A high-level summary of a role’s behavior.
action equals “dialogue”. if it’s no dialogue, summarize role’s behavior in sentence
@@
Do NOT add any clarifying information. Output MUST follow the schema above.
Do NOT add any additional columns that do not appear in the schema.
@@
You are an expert in scene analysis, skilled at analyzing and extracting key information from texts.
Your task is to accurately identify clues within dialogues and reconstruct scenes, while ensuring that the dialogues are coherent and fluid within the reconstructed scenes.
@@
Now, based on the above requirements, reconstruct the sub-scene where the dialogue takes place, and describe it accordingly.
Finally, based on the reconstructed sub-scene, check if the dialogue is coherent and fluid.
You may briefly analyze the scene context and its coherence, then return your evaluation result in a JSON-parsable format as follows:
{“scene”: “reconstructed scene description”, “coherence”: 1/0}
Where “coherence” of 1 indicates the dialogue is coherent and fluid with the scene, and 0 indicates the dialogue is not coherent with the scene.
Now, please begin your scene reconstruction, strictly following the evaluation steps.
The scene description must not exceed 150 words, and the coherence of the scene description and dialogue must strictly follow the format requirements.
@@
You are an expert in the fields of literature and psychology, skilled at analyzing and interpreting the role traits and dialogue performances in texts.
Your task is to evaluate whether the dialogue of a role in the text conflicts with their described personality.
@@
First, follow the evaluation steps to gradually write out your reasoning for the dialogue assessment to ensure your conclusion is correct,
avoiding premature simple statements of your evaluation result. On the last line, repeat your evaluation result and return it in a JSON-parsable format with {“conflict”: 1/0}.
@@
Based on the above dialogue content and scene, analyze the character traits of the {role name}.
Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the
objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
@@
{“character”: “trait1, trait2…”}
Now, please begin your analysis of {role name}’s character.
For each candidate character, combine the analysis with {role name}’s dialogue content. Finally, select the character traits from the [Candidate Character Set]
@@
Based on the dialogue content and scene above, analyze the speaking style of the {role name}.
Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal
biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
@@
Now, please begin your analysis of {role name}’s speaking style. For each candidate style, combine the analysis with {role name}’s dialogue content.
Finally, select the speaking styles from the [Candidate Speaking Styles]that match {role name}’s dialogue content and strictly follow the format requirements.
@@
Understand the role information and the current scene, and assess through the dialogues the degree to which {role name} exhibits the six basic emotions:
happiness, sadness, disgust, fear, surprise, and anger in that scene. Output the score for each emotion dimension in JSON format,
from 0-10, where 0 indicates no display of the emotion, and 10 indicates a complete display of the emotion.
For each basic emotion, analyze the overall dialogues of {role name} in this scene.
Ensure your analysis is based on the overall dialogue content and scene,
avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis,
and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
Finally, return your evaluation results in a JSON-parsable format as follows:
@@
You are an emotional analysis expert, proficient in emotional analysis, psychology, dialogue understanding, and interpersonal relationship assessment.
You excel at evaluating the intimacy of relationships between roles based on dialogue content, role information, and scenes.
@@
Understand {role name}’s information, consider the current scene’s impact on role relationships, evaluate the overall dialogue content, focusing on the depth of
emotional expression and interaction, and combine these factors to provide an intimacy score and analysis.
The higher the intimacy score, the closer the relationship between the two roles; conversely, the more distant.
The intimacy score ranges from 0-10, where 0 represents the most distant relationships, indicating strangers, hostility, indifference, etc.,
and 10 represents the closest relationships, such as lovers, kin, or friends.
Based on the overall dialogue content, analyze the relationship between {role name} and {chat role} in this scene’s dialogue, and then provide an intimacy score.
Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure
the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
Finally, return your evaluation result in a JSON-parsable format as follows:
@@
The 8 letters of the MBTI correspond as follows:
Introverted (I) / Extraverted (E); Intuitive (N) / Sensing (S); Thinking (T) / Feeling (F); Judging (J) / Perceiving (P).
@@
Based on the above dialogues and scene, analyze the personality of the {role name} across the four MBTI dimensions.
Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and
accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
Finally, return your evaluation result in a JSON-parsable format as follows:
@@
You need to design a new role description that will converse with {role name},
ensuring that the dialogue with this role effectively reflects {role name}’s personality, character traits, and speaking style.
@@
You need to creatively construct a new role setting to dialogue with {role name}, based on the traits of {role name}.
This new role should not appear in any works related to {role name}.
The description of the new role should include the role’s name and a brief personal description, to be output in JSON format like:
@@
You need to construct an engaging scene based on the information of roles A and B.
The scene and roles’ actions must be consistent with the settings of both roles, ideally within 50-100 words,
and consistent with the world they inhabit. The output should be in JSON format, like
@@
Now, please create a scene that fits the settings of both roles and is engaging.
The scene should not directly include roles’ dialogues. Ensure that the output format meets the specified requirements.
@@
You need to assign six basic emotions to the {role name} in a specific scene:
happiness, sadness, disgust, fear, surprise, and anger, based on the dialogue role’s information and the scene of the dialogue.
@@
Understand the role descriptions and the current scene, and assign the six basic emotions reflected in {role name}’s statements in that scene:
happiness, sadness, disgust, fear, surprise, and anger. Output the score for each emotion dimension in JSON format,
from 0-10, where 0 means the emotion is not displayed at all, and 10 means the emotion is fully displayed.
Please analyze in a few brief sentences the scores for the six basic emotions that {role name} should exhibit in this scene,
avoiding a simplistic statement of your evaluation results initially to ensure your conclusions are correct.
Finally, return your evaluation results in a JSON-parsable format as follows:
@@
Now, strictly follow the requirements to analyze the scene and role information,
and assign emotional scores to {role name} that must be consistent with the role’s settings and the current scene.
The analysis should be brief, not too lengthy, and avoid additional content. The final emotion scores must strictly follow the format requirements.
@@
You are an emotional analysis expert, proficient in psychology and interpersonal relationship assessment,
skilled at initializing the intimacy level of relationships based on roles’ personalities and scenarios.
You need to analyze the intimacy level between the {role name} and {chat role} by analyzing the information of both roles and the content of the scene.
@@
Understand {role name}’s personality and consider the current scene’s impact on the roles’ relationships, such as environment and context.
Integrate the above factors to initialize their intimacy score. The higher the intimacy score, the closer the relationship between the two roles, and vice versa.
The intimacy score ranges from 0-10, where 0 represents the most distant relationships,
which can indicate strangers, hostility, indifference, etc., and 10 represents the closest relationships, which can include lovers, family, and friends.
Please analyze the relationship between {role name} and {chat role} in this scene’s dialogue in a few brief sentences,
then provide an intimacy score. Avoid simply stating your evaluation results initially to ensure your conclusions are correct.
Finally, return your evaluation result in a JSON-parsable format as follows: